# Equilibrium under Time-Inconsistency: A New Existence Theory by Vanishing Entropy Regularization

### 11 Mar 2026 v1

@@ 3 Time-inconsistent Control under Entropy Regularization
Notice that $p=\frac{d+2}{1-\alpha}$. For an arbitrary fixed $(t_{0},x_{0})\in\mathbb{T}\times\mathbb{R}^{d}$, Sobolev embedding (see, e.g., [19, II.3]) gives that
$\|V^{\pi}_{\lambda}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(D_{1}(t_{0},x_{0}))}\leq A_{0}\|V^{\pi}_{\lambda}\|_{W^{1,2}_{p}(D_{2}(t_{0},x_{0}))}.$ (3.25)
$\|V^{\pi}_{\lambda}\|_{W^{1,2}(D_{2}(t_{0},x_{0}))}$ $\leq A_{0}\bigg(\|r^{\pi}\|_{L^{\infty}(D_{3}(t_{0},x_{0}))}+\|\lambda\delta\mathcal{H}(\pi)\|_{L^{\infty}(D_{3}(t_{0},x_{0}))}\bigg).$ (3.26)
This, together with (3.20) and (3.25), leads to
$\|V^{\pi}_{\lambda}\|_{W^{1,2}_{p}(D_{1}(t_{0},x_{0}))}\vee\|V^{\pi}_{\lambda}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(D_{1}(t_{0},x_{0}))}$
$\leq A_{0}\left(1+\ln\left(1+\|D_{x}w\|_{{\mathcal{C}}^{0}(\mathbb{R}^{d})}\right)\right)\quad\text{uniformly on $t_{0},x_{0}$},$
and hence,
$\|V^{\pi}_{\lambda}\|_{W^{1,2,\operatorname{\textbf{ul}}}_{p}(\mathbb{T}\times\mathbb{R}^{d})}\vee\|V^{\pi}_{\lambda}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}\leq A_{0}\left(1+\ln\left(1+\|w\|_{{\mathcal{C}}^{1}(\mathbb{R}^{d})}\right)\right).$ (3.27)
Then, it follows from (3.24) and (3.27) that,
$\|V^{\pi}_{\lambda}\|_{W^{1,2,\operatorname{\textbf{ul}}}_{p}(\mathbb{T}\times\mathbb{R}^{d})}\vee\|V^{\pi}_{\lambda}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}\vee\sup_{x\in\mathbb{R}^{d}}\|V^{\pi}_{\lambda}(\cdot,x)\|_{{\mathcal{C}}^{1}_{\alpha/2}(\mathbb{T})}$ (3.28)
$\leq A_{0}\left(1+\ln\left(1+\|w\|_{{\mathcal{C}}^{1}(\mathbb{R}^{d})}\right)\right).$
The right-hand side of (3.28) is a function on $\|w\|_{{\mathcal{C}}^{1}(\mathbb{R}^{d})}$ of sub-linear growth. We can thus obtain the existence of a constant $A^{*}$ such that for any function $w(x)$ with $\|w(x)\|_{{\mathcal{C}}^{1}_{\alpha}(\mathbb{R}^{d})}\leq A^{*}$, $V^{\pi}$ with $\pi=\Gamma_{\lambda}(x,D_{x}w(x),a)$ satisfies
$|V^{\pi}_{\lambda}\|_{W^{1,2,\operatorname{\textbf{ul}}}_{p}(\mathbb{T}\times\mathbb{R}^{d})}\vee\|V^{\pi}_{\lambda}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}\vee\sup_{x\in\mathbb{R}^{d}}\|V^{\pi}_{\lambda}(\cdot,x)\|_{{\mathcal{C}}^{1}_{\alpha/2}(\mathbb{T})}\leq A^{*}.$
@@
where $\tilde{C}$ is a constant only depending on $\|b^{\pi}\|_{{\mathcal{C}}_{\alpha}(\mathbb{R}^{d})}$ and the holder continuity of $\sigma$ such that $\tilde{C}$ can be bounded by $\varphi(\|b^{\pi}\|_{{\mathcal{C}}_{\alpha}(\mathbb{R}^{d})})$ with $\varphi(\cdot):\mathbb{R}_{+}\to\mathbb{R}_{+}$ being fixed increasing function. Therefore, by (3.2) and Lemma 3.2, it holds that
$\|V^{\pi}_{\lambda}\|_{{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(D_{1}(t_{0},x_{0}))}\leq$ $\varphi\left(\frac{A_{0}A^{*}}{\lambda}\exp(\frac{A_{0}A^{*}}{\lambda})\right)\frac{A_{0}A^{*}}{\lambda}\exp(\frac{A_{0}A^{*}}{\lambda})=:A_{\lambda},\quad\forall(t_{0},x_{0})\in\mathbb{T}\times\mathbb{R}^{d},$
@@
Next, we shall use the fixed point argument to show the existence of a regularized equilibrium. To begin with, for any $w\in{\mathcal{C}}^{0,1}(\mathbb{T}\times\mathbb{R}^{d})$, let us define $\Phi_{\lambda}(w)$ by
$\Phi_{\lambda}(w):=V^{\Gamma_{\lambda}(D_{x}w(0,x))}$ (3.29)
@@
As $u^{i+1}=u^{i}\text{ on }D_{i},$ $w$ is well-defined that satisfies
@@
$\sum_{N=N_{0}+1}^{\infty}\frac{(5N)^{\alpha-\beta}A_{\lambda}}{2^{N}}\leq\frac{\varepsilon}{4},\text{ and }\sum_{N=1}^{N_{0}}\frac{1}{2^{N}}\|w^{n,n}-w\|_{{\mathcal{C}}^{1,2}_{\beta/2,\beta}(D_{N_{0}})}\leq\frac{\varepsilon}{2}\text{ for all $n\geq n_{0}$},$
@@
$\leq\sum_{N=1}^{N_{0}}\frac{1}{2^{N}}\|w^{n,n}-w\|_{{\mathcal{C}}^{1,2}_{\beta/2,\beta}(D_{N_{0}})}+\sum_{N=N_{0}+1}^{\infty}(5N)^{\alpha-\beta}\left(\|w^{n,n}\|_{{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(D_{N})}+\|w\|_{{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(D_{N})}\right)$
@@
In conclusion, $w$ is the limit of $(w^{n,n})_{n\in\mathbb{N}}$ within ${\mathcal{M}}_{\lambda}$, and hence ${\mathcal{M}}_{\lambda}$ is a compact of the space ${\mathcal{C}}^{1,2,\operatorname{\textbf{wg}}}_{\beta/2,\beta}(\mathbb{T}\times\mathbb{R}^{d})$.
Part (ii). Take $w\in{\mathcal{M}}_{\lambda}$ such that $\|w(0,\cdot)\|_{{\mathcal{C}}^{1}_{\alpha}(\mathbb{R}^{d})}\leq\|w\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}\leq A^{*}$, Lemma 3.3 shows that $\Phi_{\lambda}(w)\in{\mathcal{M}}_{\lambda}$. Thus, $\Phi_{\lambda}$ maps from ${\mathcal{M}}_{\lambda}$ to ${\mathcal{M}}_{\lambda}$.
We next show that $\Phi_{\lambda}$ is a continuous mapping from ${\mathcal{M}}_{\lambda}$ to ${\mathcal{M}}_{\lambda}$, equipped with norm $\|\cdot\|_{{\mathcal{C}}^{1,2,\operatorname{\textbf{wg}}}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}$. Then the desired result holds due to $\beta<\alpha$.
Step 1. Take a sequence $(w^{n})_{n\in\mathbb{N}\cup\{\infty\}}\subset{\mathcal{M}}_{\lambda}$ with $\lim_{n\to\infty}\|w^{n}-w^{\infty}\|_{{\mathcal{C}}^{1,2,\operatorname{\textbf{wg}}}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}\to 0$. Set $V^{n}:=\Phi_{\lambda}(w^{n}),\pi^{n}(x,a)=\Gamma_{\lambda}(x,D_{x}w^{n}(0,x),a)$ for each $n\in\mathbb{N}\cup\{\infty\}$. By Lemma 3.1, it holds that
@@
We claim that $\|\bar{V}^{n}\|_{{\mathcal{C}}^{1,2,\operatorname{\textbf{wg}}}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}\to 0$.
Given any $\varepsilon>0$, there exists some $N_{0}$ such that $\sum_{N=N_{0}}^{\infty}\frac{A_{\lambda}}{2^{N}}\leq\frac{\varepsilon}{4}$, then $(V^{n})_{n\in\mathbb{N}\cup\{\infty\}}\in{\mathcal{M}}_{\lambda}$ gives that
$\sum_{N=N_{0}}^{\infty}\frac{1}{2^{N}}\|\bar{V}^{n}\|_{{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(D_{N})}\leq\sum_{N=N_{0}}^{\infty}\frac{1}{2^{N}}\left(\|V^{n}\|_{{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(D_{N})}+\|V^{\infty}\|_{{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(D_{N})}\right)\leq 2\sum_{N=N_{0}}^{\infty}\frac{A_{\lambda}}{2^{N}}\leq\frac{\varepsilon}{2}.$
@@
$\lim_{n\to\infty}\|\bar{V}^{n}\|_{{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(D_{N_{0}+1})}=0,$ (3.35)
Step 2. To verify (3.35), we first show
In the following proof, let $A_{0}$ be a generic finite constant depending on $A^{*},A_{\lambda}$, $N_{0},\lambda$, $\operatorname{\texttt{Leb}}(U)$ but independent of $n$, which may change from line to line. Write $g^{n}(x,a):=\frac{1}{\lambda}[b(x,a)D_{x}w^{n}(0,x,a)+r(0,x,a)]$ for $n\in\mathbb{N}\cup\{\infty\}$. Note that
$\sup_{a\in U}\|g^{n}(\cdot,a)\|_{{\mathcal{C}}_{\alpha}(B_{N_{0}+1}(0))}\leq A_{0},\sup_{a\in U}\|\exp(g^{n}(\cdot,a))\|_{{\mathcal{C}}_{\alpha}(B_{N_{0}+1}(0))}\leq A_{0},\quad\forall n\in\mathbb{N}\cup\{\infty\}.$ (3.37)
$fg]_{{\mathcal{C}}_{\alpha}(D)}\leq\|f\|_{L^{\infty}(D)}[g]_{{\mathcal{C}}_{\alpha}(D)}+\|g\|_{L^{\infty}(D)}[f]_{{\mathcal{C}}_{\alpha}(D)},\quad\left[\int_{U}h(x,a)da\right]_{{\mathcal{C}}_{\alpha}(D)}\leq\sup_{a\in U}[h(\cdot,a)]_{{\mathcal{C}}_{\alpha}(D)}.$ (3.38)
First, (3.37) and (3.38) together imply that
Direct computations yield that
$=\ln\left(\int_{U}\frac{\exp[g^{n}(x,a)-g^{\infty}(x,a)]\exp[g^{\infty}(x,a)]}{\int_{U}\exp[g^{\infty}(x,a^{\prime})]da^{\prime}}da\right)=\ln\left(\int_{U}\exp[g^{n}(x,a)-g^{\infty}(x,a)]\pi^{\infty}(x,a)da\right).$
In view that
applying (3.38) to the above equality, together with (3.37), yields
which further implies that
$\|\exp[g^{n}(x,a)]-\exp[g^{\infty}(x,a)]\|_{{\mathcal{C}}_{\alpha}(B_{N_{0}+1}(0))}\leq A_{0}\|D_{x}(w^{n}-w^{\infty})(0,\cdot)\|_{B_{N_{0}+1}(0)}.$ (3.41)
Note that
$\pi^{n}(x,a)-\pi^{\infty}(x,a)$
Applying (3.38) to two terms respectively in the last line above, together with (3.37) and (3.41), we get that
Then combining (3.42) with Assumptions 2.2 and (2.1) and (3.38) together shows
Meanwhile, it holds that
$\mathcal{H}(\pi^{n}(x))-\mathcal{H}(\pi^{\infty}(x))=$ $\int_{U}[g^{n}(x,a)-g^{\infty}(x,a)]\pi^{n}(x,a)da+\int_{U}g^{\infty}(x,a)[\pi^{n}(x,a)-\pi^{\infty}(x,a)]da$
Consequently, applying (3.37) and (3.38) to the first two integrals on the right-hand side of the equality above, together with (3.40), gives that
which, together with (3.2), yields
$\|\lambda\delta(\mathcal{H}(\pi^{n})-\mathcal{H}(\pi^{\infty}))\|_{{\mathcal{C}}_{\alpha/2,\alpha}(D_{N_{0}+1})}\leq A_{0}\|D_{x}(w^{n}-w^{\infty})\|_{{\mathcal{C}}_{\alpha}(B_{N_{0}+1}(0))}.$ (3.44)
Plugging (3.43) and (3.44) into (3.34), we readily arrive at (3.36).
Step 3. We now show that
$\|\bar{V}^{n}\|_{{\mathcal{C}}^{0,0}(D_{N_{0}+1})}\to 0.$ (3.45)
Define $\rho^{n}_{N}:=\inf\{s\geq 0:X^{\pi^{n}}_{s}\notin B_{N}(0)\}$ for $n\in\mathbb{N}\cup\{\infty\}$ and note that $\sup_{n\in\mathbb{N}\cup\{\infty\}}\|b^{\pi^{n}}\|_{{\mathcal{C}}_{\alpha}(\mathbb{R}^{d})}$ is bounded. This, together with the condition on $\sigma$ in Assumption 2.2, ensures that, for any $\varepsilon>0$, there exists some $N_{1}\in\mathbb{N}$ sufficiently large such that
$\|V^{n}(t,x)-V^{n}_{N_{1}}(t,x)\|_{{\mathcal{C}}^{0,0}(D_{N_{0}+1})}\leq\varepsilon\quad\forall n\in\mathbb{N}\cup\{\infty\}.$ (3.46)
@@
$V^{n}_{N_{1}}(t,x):=E_{x}\left[\int_{0}^{\rho^{n}_{N_{1}}\wedge(N_{1}-t)}\left(r^{\pi^{n}}(t+s,X^{\pi^{n}}_{s})+\delta(t+s)\lambda\mathcal{H}(\pi^{n}(X^{\pi^{n}}_{s}))\right)ds\right],\text{ for }(t,x)\in D_{N_{1}}.$
@@
A similar argument as in Step 2 yields that $\|r^{\pi^{n}}-r^{\pi^{\infty}}\|_{{\mathcal{C}}_{\alpha/2,\alpha}(D_{N_{1}})}$, $\|\delta\mathcal{H}({\pi^{n}})-\delta\mathcal{H}({\pi^{\infty}})\|_{{\mathcal{C}}_{\alpha/2,\alpha}(D_{N_{1}})}$, and $\|b^{\pi^{n}}-b^{\pi^{\infty}}\|_{{\mathcal{C}}_{\alpha}(B_{N_{1}}(0))}$ all tend to zero as $n\to\infty$. As a result, the term (I) tends to zero as $n\to\infty$ uniformly for $(t,x)\in D_{N_{1}}$. Also, by standard stability theory of SDE, we have
[…]
Consequently, the term (II) also tends to zero uniformly for $(t,x)\in D_{N_{1}}$. In sum, we conclude that
$\sup_{(t,x)\in D_{N_{1}}}|V^{n}_{N_{1}}(t,x)-V^{\infty}_{N_{1}}(t,x)|\to 0\text{ as $n\to\infty$}.$

### 04 Apr 2026 v2

@@ 3 Existence of Equilibrium under Entropy Regularization
By Steps 1 and 2,
$\|V^{\pi}_{\lambda}\|_{W^{1,2,\operatorname{\textbf{ul}}}_{p}([t,\infty)\times\mathbb{R}^{d})}\vee\|V^{\pi}_{\lambda}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}([t,\infty)\times\mathbb{R}^{d}))}\leq A_{0}\Psi(t)\left(1+\ln\left(1+\|w\|_{{\mathcal{C}}^{1}(\mathbb{R}^{d})}\right)\right).$ (3.31)
Observe that the function $y\mapsto A_{0}\left(1+\ln\left(1+y\right)\right)$, mapping from $0,\infty)$ to $0,\infty)$, has sub-linear growth. Consequently, there exists a constant $A^{*}$ such that
$0\leq A_{0}\left(1+\ln\left(1+y\right)\right)\leq A^{*},\text{ whenever }0\leq y\leq A^{*}.$
We can thus conclude from (3.30) and (3.31) that
$\begin{cases}\|V^{\pi}_{\lambda}\|_{W^{1,2,\operatorname{\textbf{ul}}}_{p}([t,\infty)\times\mathbb{R}^{d})}\vee\|V^{\pi}_{\lambda}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}([t,\infty)\times\mathbb{R}^{d}))}\leq\Psi(t)A^{*},\;\;\forall t\in\mathbb{T}\\ \sup_{x\in\mathbb{R}^{d}}\|V^{\pi}_{\lambda}(\cdot,x)\|_{{\mathcal{C}}^{1}_{\alpha/2}(\mathbb{T})}\leq A^{*}\end{cases}\text{provided that $\|w\|_{{\mathcal{C}}^{1}_{\alpha}(\mathbb{R}^{d})}\leq A^{*}$. }$
@@
where $\tilde{C}$ is a constant only depending on $\|b^{\pi}\|_{{\mathcal{C}}_{\alpha}(\mathbb{R}^{d})}$ and the Hölder continuity of $\sigma$ such that $\tilde{C}$ can be bounded by $\varphi(\|b^{\pi}\|_{{\mathcal{C}}_{\alpha}(\mathbb{R}^{d})})$ with a fixed increasing function $\varphi(\cdot):[0,\infty)\to[0,\infty)$. Therefore, by Lemma 3.2, (3.2) and (3.3), if $\|w\|_{{\mathcal{C}}^{1}_{\alpha}(\mathbb{R}^{d})}\leq A^{*}$, then it holds that
$\|V^{\pi}_{\lambda}\|_{{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(D_{1}(t_{0},x_{0}))}\leq\varphi\left(\frac{A_{0}A^{*}}{\lambda}\exp(\frac{A_{0}A^{*}}{\lambda})\right)\frac{A_{0}A^{*}}{\lambda}\exp(\frac{A_{0}A^{*}}{\lambda})=:A_{\lambda}\quad\forall(t_{0},x_{0})\in\mathbb{T}\times\mathbb{R}^{d},$
@@
Now we are ready to show Part (ii) of Theorem 3.1.
Proof for Theorem 3.1 Part (ii): .
Let $(u,\pi^{*})$ be a pair such that $u(t,x)\in{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})$ is a classical solution to the EEHJB system (3.5)–(3.6) and (3.7) holds. Lemma 3.3 implies that, with $\pi^{*}(x,a)=\Gamma_{\lambda}(x,D_{x}u(0,x),a)$, the function $V^{\pi^{*}}_{\lambda}(t,x)\in{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})$ is the unique solution to (3.6). Thus, $V^{\pi^{*}}_{\lambda}=u$. It remains to show that (3.4) holds for $\pi^{*}$. For any $x\in\mathbb{R}^{d}$ and $\varpi\in{\mathcal{P}}_{c}(U)$, since $V^{\pi^{*}}_{\lambda}(t,x)\in{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})$, by applying Itô formula as in the proof for Part (i) of Theorem 3.1, we obtain equation (3.10), and taking the limit as $\lambda\to 0$ then yeilds (3.11), which is bounded by the right-hand side of (3.5). Therefore, $\pi^{*}$ satisfies (3.4) and hence is a regularized equilibrium. ∎
Next, we shall use the fixed point argument to show the existence of a solution to the EEHJB system (3.5)–(3.6). To begin with, for any $w\in{\mathcal{C}}^{0,1}(\mathbb{T}\times\mathbb{R}^{d})$, let us define $\Phi_{\lambda}(w)$ by
$\Phi_{\lambda}(w):=V^{\Gamma_{\lambda}(D_{x}w(0,x))}.$ (3.32)
@@
Lemma 3.4.
Take a sequence $(w^{n})_{n\in\mathbb{N}\cup\{\infty\}}\subset{\mathcal{M}}_{\lambda}$ with $\lim_{n\to\infty}\|w^{n}-w^{\infty}\|_{{\mathcal{C}}^{1,2,\operatorname{\textbf{wg}}}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}=0$. Set $\pi^{n}(x,a)=\Gamma_{\lambda}(x,D_{x}w^{n}(0,x),a)$ for all $n\in\mathbb{N}\cup\{\infty\}$. Then for an each $N\in\mathbb{N}$, it holds that
$\lim_{n\to\infty}\left(\|r^{\pi^{n}}-r^{\pi^{\infty}}\|_{{\mathcal{C}}_{\alpha/2,\alpha}(D_{N})}+\|b^{\pi^{n}}-b^{\pi^{\infty}}\|_{{\mathcal{C}}_{\alpha}(B_{N}(0))}+\|\lambda\delta(\mathcal{H}(\pi^{n})-\mathcal{H}(\pi^{\infty}))\|_{{\mathcal{C}}_{\alpha/2,\alpha}(D_{N})}\right)=0.$
Proof.
Through the proof, let $A_{0}$ be a generic finite constant depending on $A^{*},A_{\lambda}$, $N_{0},\lambda$, $\operatorname{\texttt{Leb}}(U)$ but independent of $n$, which may change from line to line. Write $g^{n}(x,a):=\frac{1}{\lambda}[b(x,a)D_{x}w^{n}(0,x,a)+r(0,x,a)]$ for $n\in\mathbb{N}\cup\{\infty\}$. Note that
$\sup_{a\in U}\|g^{n}(\cdot,a)\|_{{\mathcal{C}}_{\alpha}(B_{N_{0}+1}(0))}\leq A_{0},\quad\sup_{a\in U}\|\exp(g^{n}(\cdot,a))\|_{{\mathcal{C}}_{\alpha}(B_{N_{0}+1}(0))}\leq A_{0},\quad\forall n\in\mathbb{N}\cup\{\infty\}.$ (3.34)
$fg]_{{\mathcal{C}}_{\alpha}(D)}\leq\|f\|_{L^{\infty}(D)}[g]_{{\mathcal{C}}_{\alpha}(D)}+\|g\|_{L^{\infty}(D)}[f]_{{\mathcal{C}}_{\alpha}(D)},\;\left[\int_{U}h(x,a)da\right]_{{\mathcal{C}}_{\alpha}(D)}\leq\sup_{a\in U}[h(\cdot,a)]_{{\mathcal{C}}_{\alpha}(D)}.$ (3.35)
First, (3.34) and (3.35) together imply that
Direct computations yield that
$=\ln\left(\int_{U}\frac{\exp[g^{n}(x,a)-g^{\infty}(x,a)]\exp[g^{\infty}(x,a)]}{\int_{U}\exp[g^{\infty}(x,a^{\prime})]da^{\prime}}da\right)$
$=\ln\left(\int_{U}\exp[g^{n}(x,a)-g^{\infty}(x,a)]\pi^{\infty}(x,a)da\right).$
In view that
applying (3.35) to the above equality, together with (3.34), yields
which further implies that
$\|\exp[g^{n}(x,a)]-\exp[g^{\infty}(x,a)]\|_{{\mathcal{C}}_{\alpha}(B_{N_{0}+1}(0))}\leq A_{0}\|D_{x}(w^{n}-w^{\infty})(0,\cdot)\|_{{\mathcal{C}}_{\alpha}(B_{N_{0}+1}(0))}.$ (3.38)
Note that
$\pi^{n}(x,a)-\pi^{\infty}(x,a)$
Applying (3.35) to the two terms respectively in the last line above, together with (3.34) and (3.38), we get that
Then combining (3.39) with Assumptions 2.2, (2.1) and (3.35) yields
Meanwhile, it holds that
$\mathcal{H}(\pi^{n}(x))-\mathcal{H}(\pi^{\infty}(x))=$ $\int_{U}g^{\infty}(x,a)[\pi^{\infty}(x,a)-\pi^{n}(x,a)]da+\int_{U}[g^{\infty}(x,a)-g^{n}(x,a)]\pi^{n}(x,a)da$
Applying (3.34) and (3.35) to the first two integrals on the right-hand side, together with (3.37), gives that
which, together with (3.2), yields
$\|\lambda\delta(\mathcal{H}(\pi^{n})-\mathcal{H}(\pi^{\infty}))\|_{{\mathcal{C}}_{\alpha/2,\alpha}(D_{N_{0}+1})}\leq A_{0}\|D_{x}(w^{n}-w^{\infty})\|_{{\mathcal{C}}_{\alpha}(B_{N_{0}+1}(0))}\to 0.$
@@
As $u^{i+1}=u^{i}\text{ on }D_{i},$ $w$ is well-defined and satisfies that
@@
$\sum_{N=N_{0}+1}^{\infty}\frac{A_{\lambda}}{2^{N}}\leq\frac{\varepsilon}{4},\;\text{ and }\sum_{N=1}^{N_{0}}\frac{1}{2^{N}}\|w^{n,n}-w\|_{{\mathcal{C}}^{1,2}_{\beta/2,\beta}(D_{N_{0}})}\leq\frac{\varepsilon}{2}\;\text{ $\forall n\geq n_{0}$},$
@@
$\leq\sum_{N=1}^{N_{0}}\frac{1}{2^{N}}\|w^{n,n}-w\|_{{\mathcal{C}}^{1,2}_{\beta/2,\beta}(D_{N_{0}})}+\sum_{N=N_{0}+1}^{\infty}\left(\|w^{n,n}\|_{{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(D_{N})}+\|w\|_{{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(D_{N})}\right)$
@@
In conclusion, $w$ is the limit of $(w^{n,n})_{n\in\mathbb{N}}$ within ${\mathcal{M}}_{\lambda}$, and hence ${\mathcal{M}}_{\lambda}$ is a compact subset of the space ${\mathcal{C}}^{1,2,\operatorname{\textbf{wg}}}_{\beta/2,\beta}(\mathbb{T}\times\mathbb{R}^{d})$.
Part (ii). Take an arbitrary function $w\in{\mathcal{M}}_{\lambda}$, we have that $\|w(0,\cdot)\|_{{\mathcal{C}}^{1}_{\alpha}(\mathbb{R}^{d})}\leq\|w\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}\leq A^{*}$, then Lemma 3.3 shows that $\Phi_{\lambda}(w)\in{\mathcal{M}}_{\lambda}$. Consequently, $\Phi_{\lambda}$ maps from ${\mathcal{M}}_{\lambda}$ to ${\mathcal{M}}_{\lambda}$. We next show that $\Phi_{\lambda}$ is a continuous mapping from ${\mathcal{M}}_{\lambda}$ to ${\mathcal{M}}_{\lambda}$, equipped with norm $\|\cdot\|_{{\mathcal{C}}^{1,2,\operatorname{\textbf{wg}}}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}$. Then the desired result holds due to $\beta<\alpha$.
Take a sequence $(w^{n})_{n\in\mathbb{N}\cup\{\infty\}}\subset{\mathcal{M}}_{\lambda}$ with $\lim_{n\to\infty}\|w^{n}-w^{\infty}\|_{{\mathcal{C}}^{1,2,\operatorname{\textbf{wg}}}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}\to 0$. Set $V^{n}:=\Phi_{\lambda}(w^{n}),\pi^{n}(x,a)=\Gamma_{\lambda}(x,D_{x}w^{n}(0,x),a)$ for each $n\in\mathbb{N}\cup\{\infty\}$. By Lemma 3.1, it holds that
@@
We now claim that $\|\bar{V}^{n}\|_{{\mathcal{C}}^{1,2,\operatorname{\textbf{wg}}}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}\to 0$. Given any $\varepsilon>0$, there exists some $N_{0}$ such that $\sum_{N=N_{0}+1}^{\infty}\frac{A_{\lambda}}{2^{N}}\leq\frac{\varepsilon}{4}$, then $(V^{n})_{n\in\mathbb{N}\cup\{\infty\}}\in{\mathcal{M}}_{\lambda}$ gives that
$\sum_{N=N_{0}+1}^{\infty}\frac{1}{2^{N}}\|\bar{V}^{n}\|_{{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(D_{N})}\leq\sum_{N=N_{0}+1}^{\infty}\frac{1}{2^{N}}\left(\|V^{n}\|_{{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(D_{N})}+\|V^{\infty}\|_{{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(D_{N})}\right)\leq 2\sum_{N=N_{0}+1}^{\infty}\frac{A_{\lambda}}{2^{N}}\leq\frac{\varepsilon}{2}.$
@@
$\lim_{n\to\infty}\|\bar{V}^{n}\|_{{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(D_{N_{0}})}=0.$ (3.45)
We begin by showing that
$\lim_{n\to\infty}\|\bar{V}^{n}\|_{L^{\infty}(D_{N_{0}+1})}=0.$ (3.46)
Define $\rho^{n}_{N}:=\inf\{s\geq 0:X^{\pi^{n}}_{s}\notin B_{N}(0)\}$ for $n\in\mathbb{N}\cup\{\infty\}$ and note that $\sup_{n\in\mathbb{N}\cup\{\infty\}}\|b^{\pi^{n}}\|_{{\mathcal{C}}_{\alpha}(\mathbb{R}^{d})}$ is bounded. This, together with the condition on $\sigma$ in Assumption 2.2 and the uniform bound on $\|V^{n}\|_{L^{\infty}(\mathbb{T}\times\mathbb{R}^{d})}$, ensures that, for any $\varepsilon>0$, there exists a sufficiently large $N_{1}\in\mathbb{N}$ with $N_{1}\geq N_{0}+1$ such that
$\|V^{n}(t,x)-V^{n}_{N_{1}}(t,x)\|_{L^{\infty}(D_{N_{0}+1})}\leq\varepsilon\quad\forall n\in\mathbb{N}\cup\{\infty\}.$ (3.47)
@@
$V^{n}_{N_{1}}(t,x):=E_{x}\left[\int_{0}^{\rho^{n}_{N_{1}}\wedge(N_{1}-t)}\left(r^{\pi^{n}}(t+s,X^{\pi^{n}}_{s})+\delta(t+s)\lambda\mathcal{H}(\pi^{n}(X^{\pi^{n}}_{s}))\right)ds\right],\text{ for }(t,x)\in D_{N_{0}+1}.$
@@
By Lemma 3.4, $\|r^{\pi^{n}}-r^{\pi^{\infty}}\|_{{\mathcal{C}}_{\alpha/2,\alpha}(D_{N_{1}})}$, $\|\delta\mathcal{H}({\pi^{n}})-\delta\mathcal{H}({\pi^{\infty}})\|_{{\mathcal{C}}_{\alpha/2,\alpha}(D_{N_{1}})}$ and $\|b^{\pi^{n}}-b^{\pi^{\infty}}\|_{{\mathcal{C}}_{\alpha}(B_{N_{1}}(0))}$ all tend to zero as $n\to\infty$. As a result, the term (I) tends to zero as $n\to\infty$ uniformly for $(t,x)\in D_{N_{0}+1}$. Also, by taking $X^{\pi^{n}}_{0}=x$ for all $n\in\mathbb{N}\cup\{\infty\}$, the standard stability theory of SDEs yields the following convergences hold uniformly over the starting position $x\in B_{N_{0}+1}(0)$:
$\sup_{0\leq s\leq N_{1}}|X^{\pi^{n}}-X^{\pi^{\infty}}|\to 0\text{ in probability },\rho^{n}_{N_{1}}\to\rho^{\infty}_{N_{1}}\text{ in probability}.$
Consequently, the term (II) also tends to zero uniformly for $(t,x)\in D_{N_{0}+1}$. In sum, we conclude that
$\sup_{(t,x)\in D_{N_{0}+1}}|V^{n}_{N_{1}}(t,x)-V^{\infty}_{N_{1}}(t,x)|\to 0\text{ as $n\to\infty$}.$
@@
By Lemma 3.4 and that $\sup_{n\in\mathbb{N}}\|V^{n}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha/2}(\mathbb{T}\times\mathbb{R}^{d})}\leq A^{*}$, $f^{n}$ defined in (3.44) satisfies that